OpenStudy (anonymous):

Does anyone know how to do this? :)) Multiply. (x + 8)(x - 6) Answers: x2 + 2x - 48 x2 - 48 x2 - 2x - 48 x2 - 48x + 2

OpenStudy (anonymous):

(x + 8)(x - 6)=x^2 - 6x + 8x - 48 = x^2 + 2x - 48 the 1st option.

OpenStudy (anonymous):

Thank you so much...you're such a help! :))